OpenStudy (anonymous):

Soybean meal is 18% protein, cornmeal is 9% protein. How many pounds of each should be mixed together in order to get a 360lb mixture that is 13% protein? Can someone help me with this?

OpenStudy (anonymous):

let x = amount of soybean and 360-x = amount of cornmeal in the mixture, then we have 18x + 9(360-x) = 13(360) 18x + 3240-9x = 4680 9x = 1440 x = 160 pounds of soybean 360-x = 200 pounds of cornmeal

OpenStudy (anonymous):

160lbs of soybean and 200lbs. of cornmeal. Thanks beans1992